solve showing work (x+3)(3x+5)=7
Answer by mananth(16949)

(x+3)(3x+5)=7
3x^2+5x+9x+15-7=0
3x^2+14x+8=0
3x^2+2x+12x+8=0
x(3x+2)+4(3x+2)=0
(x+4)(3x+2)=0
x=-4 OR x=-2/3